OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 610! If four times the number of one-step equations minus twice the number of two-step equations is equal to 340, how many one-step equations auditioned today? 260 340 350 670

OpenStudy (anonymous):

set up equations, have x equal to 1-step and y equal to 2-step x + y = 610 4x - 2y = 340 find a way to substitute for x or y see we have x and 4x , lets try that get x by itself for the first equation x = 610 - y 4x = 4(610 - y) 2440 - 4y - 2y = 340 use this equation to find # of 2-step equations then subtract from 610

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Your welcome! :)